[Brahms-l] Shift report 20060613 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Jun 13 2006 - 08:00:55 EDT
Shifter leader was rk
Shifters were nt


1. Machine development until 4.30 with a very unstable beam. Since then we have beam.
2. Runs 16112 - 16116 with FS @ 3deg @ 1723B and MRS @ 45deg @1050B
3. Several problems:
- Shower Max R1 tripped at the beginning of shift.
- Connection with FB3 was lost, restarting it wasn't easy, but finally we succeeded. Unfortunately we resetted low voltages for the DC - not all came back to what it used to be before. First we had no counts in several planes. Now we have them, but they seem to be a little low.
- Frame L1458A was off, we had to turn it on, and also voltages on TPCs and DVM.
_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Jun 13 08:01:38 2006

This archive was generated by hypermail 2.1.8 : Tue Jun 13 2006 - 08:01:53 EDT